# ULTHO: Ultra-Lightweight yet Efficient Hyperparameter Optimization in Deep Reinforcement Learning

### v1

@@ 5.2.4 Ablation Studies
While ULTHO achieves efficient HPO, the utilized UCB method relies on two HPs, the exploration coefficient $c$ and the length of sliding window $W$. To evaluate the robustness of ULTHO, we perform experiments using different configurations of the two HPs on the Procgen benchmark, and the aggregated training performance is illustrated in Figure 8. It is evident that the ULTHO is relatively insensitive to the choice of $c$ and $W$, yet a bigger $W$ helps achieve a more reliable return estimation. Therefore, ULTHO maintains stable performance across different HP and can be easily adapted to various training scenarios.
@@ 6 Discussion
In this paper, we investigated the problem of HPO in deep RL and proposed an ultra-lightweight yet powerful framework entitled ULTHO, which performs HPO on the fly and requires no complex learning process. ULTHO formulates the HPO process as a multi-armed bandit with clustered arms, enabling efficient and adaptive HP selection across different learning stages. We evaluate ULTHO on ALE, Procgen, MiniGrid, and PyBullet benchmarks. Extensive experiments demonstrate that ULTHO can effectively enhance RL algorithms with simple architecture, contributing to the development of advanced and automated RL systems.
@@ B.1 Arcade Learning Environment
PPO+Relay-ULTHO. At the end of the PPO+ULTHO experiments, we count the number of times each cluster is selected and find out the cluster of interest and neglected cluster of interest. Then we perform the experiments with the two clusters separately before reporting the best-performing cluster. Therefore, the actual training budget of Relay-ULTHO is three times that of ULTHO, i.e., 30M environment steps. Similarly, we run a grid search over the exploration coefficient and the sliding window size and report the best results.
@@ B.2 Procgen
PPO+ULTHO. In this part, we utilize the implementation of CleanRL for the PPO algorithm and train the agent for 25M environment steps on 200 levels before testing them on the full distribution of levels. For the hyperparameter clusters, we select the batch size, value loss coefficient, entropy loss coefficient, and number of update epochs as the candidates, and the detailed values of each cluster are listed in Table 2. Similarly, we run a grid search over the exploration coefficient $c\in\{1.0,5.0\}$ and the size of the sliding window used to compute the $Q$-values $W\in\{10,50,100\}$ to study the robustness of ULTHO. Finally, Table 3 illustrates the PPO hyperparameters, which remain fixed throughout all the experiments except for the hyperparameter clusters.
@@ Appendix C Learning Curves
Appendix D Detailed Decision Processes

### v2

@@ 5.2.4 Ablation Studies
While ULTHO achieves efficient HPO, the utilized UCB method relies on two HPs, the exploration coefficient $c$ and the length of the sliding window $W$. To evaluate the robustness of ULTHO, we perform experiments using different configurations of the two HPs on the Procgen benchmark, and the aggregated training performance is illustrated in Figure 8. It is evident that the ULTHO is relatively insensitive to the choice of $c$ and $W$, yet a bigger $W$ helps achieve a more reliable return estimation. Therefore, ULTHO maintains stable performance across different HPs and can be easily adapted to various training scenarios. More ablation studies regarding the number of HP clusters and training budgets can be found in Appendix D.
@@ 6 Discussion
In this paper, we investigated the problem of HPO in deep RL and proposed an ultra-lightweight yet powerful framework entitled ULTHO, which performs HPO on the fly and requires no complex learning process. ULTHO formulates the HPO process as a multi-armed bandit with clustered arms, enabling efficient and adaptive HP selection across different learning stages. We evaluate ULTHO on ALE, Procgen, MiniGrid, and PyBullet benchmarks. Extensive experiments demonstrate that ULTHO can effectively enhance RL algorithms with a simple architecture, contributing to the development of advanced and automated RL systems.
@@
Acknowledgements
This work is supported, in part, by the HKSAR RGC under Grant No. PolyU 15224823, the Guangdong Basic and Applied Basic Research Foundation under Grant No. 2024A1515011524, the NSFC under Grant No. 62302246, the ZJNSFC under Grant No. LQ23F010008, and the Ningbo under Grants No. 2023Z237 & 2024Z284 & 2024Z289 & 2023CX050011 & 2025Z038. We thank the high-performance computing center at Eastern Institute of Technology and Ningbo Institute of Digital Twin for providing the computing resources.
@@ B.1 Arcade Learning Environment
PPO+Relay-ULTHO. At the end of the PPO+ULTHO experiments, we count the number of times each cluster is selected and find out the cluster of interest and the neglected cluster of interest. Then we perform the experiments with the two clusters separately before reporting the best-performing cluster. Therefore, the actual training budget of Relay-ULTHO is three times that of ULTHO, i.e., 30M environment steps. Similarly, we run a grid search over the exploration coefficient and the sliding window size and report the best results.
@@ B.2 Procgen
PPO+ULTHO. In this part, we utilize the implementation of CleanRL for the PPO algorithm and train the agent for 25M environment steps on 200 levels before testing it on the full distribution of levels. For the hyperparameter clusters, we select the batch size, value loss coefficient, entropy loss coefficient, and number of update epochs as the candidates, and the detailed values of each cluster are listed in Table 2. Similarly, we run a grid search over the exploration coefficient $c\in\{1.0,5.0\}$ and the size of the sliding window used to compute the $Q$-values $W\in\{10,50,100\}$ to study the robustness of ULTHO. Finally, Table 3 illustrates the PPO hyperparameters, which remain fixed throughout all the experiments except for the hyperparameter clusters.
@@ Appendix C Learning Curves
Appendix D Ablation Studies
ULTHO versus baselines (same training budgets). The Figure below compares ULTHO and the baselines on the ALE-5 benchmark using the same training budget, 100M environment steps. By achieving highly frequent HP tuning within a limited budget, ULTHO consistently outperforms the baselines.
Figure 21: The aggregated performance comparison of ULTHO and baselines on the ALE-5 benchmark with the same training budgets.
ULTHO’s performance with varying sets of HPs. The Figure below compares the performance of ULTHO on the Procgen benchmark with the varying number of HP clusters and intra-cluster HPs. There is a notable performance gain when using more HP clusters. However, as a lightweight framework and constrained by the capability of MAB algorithms, continually increasing the types and granularity of HPs only achieves limited performance gains.
Figure 22: ULTHO’s performance comparison with the varying number of HP clusters and intra-cluster HPs.
ULTHO’s performance with off-policy RL algorithms. It is simple to apply ULTHO to off-policy RL algorithms. Take SAC [26] for example, every $T$ steps, we use ULTHO to select a HP before using it for model update in the next $T$ steps. The utility of each HP can be evaluated using the action value function: $U_{i}(\bm{\psi})=\frac{1}{W}\sum_{j=1}^{W}\frac{1}{|\mathcal{B}_{j}|}\sum_{(\bm{s},\bm{a})\sim\mathcal{B}_{j}}Q(\bm{s},\bm{a})$, where $\mathcal{B}$ is a sampled batch. We perform an experiment using two tasks from the DMC benchmark [57], and the Figure above indicates that ULTHO effectively enhances SAC’s performance. However, at the current stage, we do not recommend applying ULTHO to off-policy algorithms as the return estimation based on sampling from a replay buffer is relatively unstable.
Figure 23: The performance of SAC+ULTHO on the DMC benchmark.
Appendix E Detailed Decision Processes